OpenStudy (anonymous):

I need help with 4 questions please help??

OpenStudy (anonymous):

Is this a test?

OpenStudy (anonymous):

No its extra work that she wants us to do to help us understand

OpenStudy (anonymous):

Okay what are the questions?

OpenStudy (anonymous):

okay what are they?? and what grade?

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

SOlve each compound inequality and graph the solution set -4 less than or equal to -2x + 6 less than or equal to 13

OpenStudy (anonymous):

my mind is blank :) sorry bout that maybe @BlackVeilCupcakes can help you

OpenStudy (anonymous):

Solve each absolute value inequality and graph the solution set -4 |3z +5| less than or equal to 20

OpenStudy (anonymous):

@starlightstarbright thanks

OpenStudy (anonymous):

Hmm, yeah i'm not good with these things. Lol have you tried looking for a walkthrough on google about how to do equations like this?

OpenStudy (anonymous):

no prob, hehe

OpenStudy (anonymous):

I have I have but its not really helping

OpenStudy (anonymous):

-4 <= -2x + 6 <= 13 --- subtract 6 -4 - 6 <= -2x <= 13 - 6 -10 <= -2x <= 7 -- divide by -2 and change the sign 5 >= x >= -7/2 ============== -4 | 3z + 5 | < = 20 -- divide by -4 and change the sign | 3z + 5 | >= - 5 3z + 5 >= -5 3z >= -5 + 5 3z >= 0 z >= 0 -(3z + 5) >= -5 -3z - 5 >= -5 -3z >= -5 + 5 -3z >= 0 z <= 0 I am not sure about this one....because I am getting z >= 0 and z <= 0 ???

OpenStudy (anonymous):

Thank you soo much so the compund inequality is -5 < X <0 ?

OpenStudy (anonymous):

you mean the first problem ?? it is : 5 >= x >= -7/2

OpenStudy (anonymous):

I didnt see the line that divided it thanks!

OpenStudy (anonymous):

do you know how to graph them both?

OpenStudy (anonymous):

I see a mistake I made on the second one..with the absolute values 3z + 5 >= -5 3z >= -5-5 3z >= -10 z >= -10/3 so the solution for this one is : z >= -10/3 and z <= 0 =============== graphing... compound inequality.. 5 >= x >= -7/2 closed circle on 5, closed circle on -7/2(or - 3 1/2)...shading in between absolute value one -10/3 <= z <= 0 closed circle on -10/3 (or - 3 1/3), closed circle on 0, shading in between